OpenStudy (anonymous):

The length of a rectangle is 1m more than twice its width, and the area of the rectangle is 66m2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

i need the length ? and the width ?

OpenStudy (anonymous):

any helping hand please

OpenStudy (anonymous):

length = 2(width) + 1 length x width = area

OpenStudy (anonymous):

let width is x then length is 2x+1 now area is given 66 m^2 area of rectangle A = length * breadth 66 = x(2x + 1) 66 = 2x^2 +x 2x^2 + x - 66 =0 x^2 +x/2 - 33 = 0 (x+6)(x-11/2)=0 x = -6 or x = 11/2 but length can not be negative so x = 11/2 length = 2x + 1 = 2*11/2 + 1 =12 width = x = 11/2